A hybrid electric vehicle is equipped with a battery with 12 kWh and 200V. If the battery is being charge at home with 15A current, how long doe it take to charge the battery?
Answer #1

We know that power of the battery

P=VI

P=200*15=3000 W=3kW ...1

Energy E=PT

T=E/P

HERE E=12kWh ,P=3kW

so

T=12/3 =4h

T=4 h

in seconds

T=14400 sec